OpenStudy (anonymous):

6x=7-4y;x=-2,-1,0. solve equation for Y. find the value of Y for each value of X.

OpenStudy (anonymous):

i really need some help :(

OpenStudy (fanduekisses):

Ok just plug in every value of x into each x in the equation, then solve for y :)

OpenStudy (fanduekisses):

Let's try pluging in -2 first:

OpenStudy (fanduekisses):

\[6*-2=7-4y\] \[-12=7-47\]

OpenStudy (anonymous):

okay

OpenStudy (fanduekisses):

Subtract 7 to both sides

OpenStudy (fanduekisses):

ooops it's supposed to be -12=7-4y NOT −12=7−47 my bad

OpenStudy (fanduekisses):

Ok what do you get after subtracting the 7 to both sides?

OpenStudy (anonymous):

that's fine i understand. how did you get 6x?

OpenStudy (fanduekisses):

from the equation you were given "6x=7-4y"

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

i saw that now.

OpenStudy (fanduekisses):

ok what does y equal when x=-2?

OpenStudy (anonymous):

um...i think y equals i need to write more down

OpenStudy (fanduekisses):

ok you want to solve for y after plugging in the value of x

OpenStudy (anonymous):

what do you mean by plugging in?

OpenStudy (fanduekisses):

Ok so you were given three values of x, x=-2,-1,0, you were asked to find the value of y for each value of x, You use the equation given is 6x=7-4y

OpenStudy (fanduekisses):

So plug in each value of x into the equation (or substitute for x)

OpenStudy (fanduekisses):

For example when x=-2 \[6 * -2 = 7-4y\]

OpenStudy (fanduekisses):

Did you see how I put the -2 in the place where the x was?

OpenStudy (anonymous):

yes i see why?

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

so thats why you replace it.

OpenStudy (fanduekisses):

Because that's how you do it. If you are given an equation, in this case 6x=7-4y, and are given the value of the x, you plug in that value.

OpenStudy (anonymous):

so do i multiply?

OpenStudy (fanduekisses):

You are asked to find the value of the y when x equals that value you were given

OpenStudy (fanduekisses):

Try it, plug in the first value of x that you were given.

OpenStudy (anonymous):

let me try.

OpenStudy (anonymous):

6x-2=7-4y now you subtract 7 from both sides and you get -2-7

OpenStudy (anonymous):

equals -5

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

or is it -9?

OpenStudy (anonymous):

:/

OpenStudy (fanduekisses):

No you were supposed to substitute the x for -2

OpenStudy (fanduekisses):

6 * -2 = 7-4y six times -2 equals -12 -12=7-4y NOW you can subtract the 7 to both sides

OpenStudy (anonymous):

oops sorry i meant to put 6-2

OpenStudy (anonymous):

you put a * in your previous message does that mean times?

OpenStudy (anonymous):

multiplication.

OpenStudy (fanduekisses):

Yes * means times

OpenStudy (anonymous):

i subtracted 7 from both sides is that correct atleast.

OpenStudy (anonymous):

okay good.

OpenStudy (anonymous):

i was supposed to put 6 times -2 =7 -4y

OpenStudy (fanduekisses):

Yes, because the 6x just meant 6 times x, and x is -2

OpenStudy (anonymous):

okay i understand.

OpenStudy (anonymous):

what do i do with 4y?

OpenStudy (anonymous):

-4y sorry.

OpenStudy (fanduekisses):

Ok so after plugging in and multiplying you have -12=7-4y, the goal here is to get y by itself, like (y= something)

OpenStudy (fanduekisses):

**Oops and after subtracting the 7

OpenStudy (anonymous):

wait i think i already know that you divide it with the answer of -2-7 right?

OpenStudy (anonymous):

that was -5 when i calculated it.

OpenStudy (fanduekisses):

so the -4y, just means -4 times y, so you do the opposite, you divide -4 to both sides

OpenStudy (anonymous):

is 2-7 negative 5 or -9?

OpenStudy (anonymous):

y=1.25

OpenStudy (anonymous):

or no :/

OpenStudy (fanduekisses):

\[-12=7-4y\] you subtract 7 to both sides \[-19=4y\] you divide 4 to both sides \[\frac{ -19 }{ 4}=y\]

OpenStudy (fanduekisses):

Where did you get the 2-7?

OpenStudy (anonymous):

i subtracted 7 from both sides i guess i had to subtact on the right side of the problem not the left.

OpenStudy (fanduekisses):

Look after plugging in the -2, you get 6 * -2=7-4y then you multiply that two you get: -12=7-4y then you subtract the 7 to both sides (the 7's cancel each other out)

OpenStudy (anonymous):

oh i get it now!

OpenStudy (anonymous):

i forgot that step.

OpenStudy (anonymous):

i canceled the 7 outs tho.

OpenStudy (anonymous):

i'll figure it out.

OpenStudy (anonymous):

thank you

OpenStudy (fanduekisses):

Look at what I did click the image

OpenStudy (anonymous):

I see i accidentally solved it wrong.

OpenStudy (anonymous):

thanks for drawing the picture i see it now what i did wrong, you subtracted the 7 from -12 thats how you got -19.

OpenStudy (anonymous):

thanks i don't need anymore help i got it.

OpenStudy (anonymous):

now i just have to find the value of y for the other x of 4